a) Does there exist a plane targent to x
2 −2y
2 +2z
2 = 8 and which passes through
2x+3y+2z = 8, x−y+2z = 5? Justify your answer. (5)
1
Expert's answer
2020-03-17T10:48:35-0400

Given "f(x,y,z)=x^{2}-2y^{2}+2z^{2}-8=0" with point "P(a,b,c)" .


We calculate "(f_{x},f_{y},f_{z})=(2x,-4y,4z)" in this point: "(2a,-4b,4c)" .


Equation of the tangent plane is "2a(x-a)-4b(x-b)+4c(x-c)=0" , "ax-2by+2cz=8."


The system "\left\{\begin{array}{ll} 2x+3y+2z = 8 \\ x-y+2z=5 \end{array}\right." is a line, we choose points on the line.


The points "(3,0,1)" and "(-1,1,\frac{7}{2})" of the given line must lie in this tangent plane,


then we have "c=\frac{8-3a}{2}" and "b=\frac{40-23a}{4}" .


Using the equation "a^{2}-2b^{2}+2c^{2}-8=0" , we do not find a real solution for "a" .


Answer: The tangent plane does not exist.